# CustomVideo: Customizing Text-to-Video Generation with Multiple Subjects

### v1

@@ Appendix B CustomStudio Dataset
We collect a dataset CustomStudio for multi-subject driven video generation. This dataset is composed of 69 individual objects and 57 meaningful pairs. The objects cover a wide range of categories, including decor item, food, furniture, instrument, jewelry, luggage, person, pet, plant, plushie, scene, thing, toy, transportation, and wearable item. The images are adopted from DreamBooth Ruiz et al. (2023), CustomDiffusion Kumari et al. (2023) and internet. An overview of the samples in CustomStudio dataset is shown in Figure 7.
Appendix C More Experimental Details
All of the training process is under fp16 mixed precision with accelerate package Gugger et al. (2022). We use data augmentations during training, including randomly horizontal flip, randomly crop and resize, together with corresponding prompt change (‘very small’ or ‘close up’ appends before the prompt).
DreamBooth.
For efficient fine-tuning, we utilize LoRA Hu et al. (2022) to adapt the UNet under DreamBooth Ruiz et al. (2023). The rank of LoRA module is set as 4. We use AdamW Loshchilov and Hutter (2017) optimizer, with learning rate 4e-4, weight decay 1e-2, batch size 2, and training steps 1000.
CustomDiffusion.
The weights of query and value in all of the cross attention layers of UNet are fine-tuned in CustomDiffusion Kumari et al. (2023). We use AdamW Loshchilov and Hutter (2017) optimizer, with learning rate 4e-5, weight decay 1e-2, batch size 2, and training steps 500.
VideoDreamer.
LoRA Hu et al. (2022) is utilized for fine-tuning the UNet and text encoder in VideoDreamer Chen et al. (2023b). The AdamW Loshchilov and Hutter (2017) optimizer is used with learning rate 5e-5, weight decay 1e-2, batch size 2, and training steps 500.
CustomVideo.
The weights of query and value in all of the cross attention layers of UNet are fine-tuned. The learnable word tokens are initialized with the tokens of the corresponding class names. The sampling step for high resolution video is 30.
Table 5: Quantitative results of cross attention levels. We find that only using cross attention map from $\ell_{3}$ works best.
@@ Appendix C More Experimental Details
0.6134 0.5759 0.3023 0.8125
0.6694 0.6239 0.3450 0.7346
0.7051 0.6749 0.3955 0.8142
0.6144 0.6545 0.3426 0.7992
0.6882 0.6177 0.3010 0.8076
0.7047 0.5801 0.3515 0.7930
0.6502 0.6407 0.3651 0.7875
0.6267 0.6734 0.3746 0.7594
0.6869 0.5885 0.3894 0.7605
Figure 7: The overview of our CustomStudio dataset. Samples in our CustomStudio dataset cover a wide range of object categories.

### v2

@@ Appendix A CustomStudio Dataset
We collect a dataset CustomStudio for multi-subject driven video generation. This dataset is composed of 63 individual objects and 68 meaningful pairs. These objects cover a wide range of 13 diverse categories, including anime character, decorate item, furniture, instrument, person, pet, plant, plush, scene, thing, toy, transport, and wearable item. For each subject, 4-21 images are provided as the reference images. The images are adopted from DreamBooth [45], CustomDiffusion [40] and Mix-of-Show [52]. An overview of the samples in CustomStudio dataset is shown in Figure 8. A meaningful pair is composed of different subjects may appear in a video together in practice. For example, ‘cat’ and ‘dog’ can be a meaningful pair, while ‘book’ and ‘barn’ can not arise in the most cases. We set up 25 types of subject pairs, which are constructed based on the corresponding categories, as shown in Table 4.
Two-subject Pair [decorate item, plant], [person, transport], [pet, pet], [furniture, thing], [toy, toy]
[decorate item, furniture], [instrument, person], [person, person], [plush, plush]
[anime character, anime character], [toy, transport], [person, wearable item]
[person, scene], [scene, transport], [instrument, anime character]
Three-subject Pair [anime character, anime character, scene], [anime character, instrument, scene]
[toy, toy, toy], [plush, plush, plush], [person, instrument, scene]
[person, person, scene], [pet, wearable item, scene], [toy, transport, scene]
[pet, pet, scene], [decorate item, furniture, furniture]
Table 4: Subject pairs constructed upon categories of the subjects in our CustomStudio dataset.
Prompt Templates.
The prompt templates used for generating videos during inference are shown in Table 5. For example, we want to generate a customized video with a ‘cat’ and a ‘dog’, the prompt for inference can be ‘a <new1> cat and a <new2> dog walking in the Times Square’.
Prompt Template for 2-subject Pair
a <c1> and a <c2> sitting on an antique table
a <c1> and a <c2> sitting on beach with a view of seashore
a <c2> and a <c1> side by side on a mountaintop, overlooking a sunrise
a <c1> and a <c2> on a surfboard together in the middle of a clear blue ocean
a <c1> playing with <c2>
a <c1> playing with a robot toy <c2>
a <c2> playing with a robot toy <c1>
a plush toy replica of a <c1> and a <c2> sitting beside it
a <c1> and a <c2> walking in the Times Square
a <c1> and a <c2> walking on the Great Wall
Prompt Template for 3-subject Pair
a <c1> wearing a <c2>, with a <c3> in the background
a <c1> wearing a <c2> and giving a speech, with a <c3> in the background
a wide shot of a <c1> wearing a <c2> with boston city in background, with a <c3> in the background
a long shot of a <c1> walking their golden retriever, wearing a <c2>, with a <c3> in the background
a <c1> sitting on the sidewalk wearing a <c2>, with a <c3> in the background
a <c1> standing at a graffiti wall, showcasing a new <c2>, with a <c3> in the background
close shot of a <c1> walking on a ramp wearing a <c2>, with a <c3> in the background
a <c1> stepping out of a taxi in a <c2>, with a <c3> in the background
close shot of a <c1> walking in rain wearing a <c2>, with a <c3> in the background
long shot of a <c1> sitting on the edge of a roof wearing <c2>, with a <c3> in the background
Table 5: Prompt templates used for generating videos during inference.
Appendix B Additional Experimental Results
Layers of Cross-attention Maps.
We investigates the effect of how we extract the cross-attention maps from the cross-attention layers. There are four levels of cross-attention layers in the U-Net, in which the sizes cross-attention maps are different. Taking the resolution $576\times 320$ of our generated low resolution video as an example, the sizes of cross-attention maps are $72\times 40$ ($\mathcal{\ell}_{1}$), $36\times 20$ ($\mathcal{\ell}_{2}$), $18\times 10$ ($\mathcal{\ell}_{3}$), and $9\times 5$ ($\mathcal{\ell}_{4}$). Here, we study the effect of different levels of cross-attention maps, the results are shown in Table 6. We observe a significant performance drop with too large ($\mathcal{\ell}_{1}$) or small ($\mathcal{\ell}_{4}$) cross-attention maps. Large cross-attention maps may, on one hand, lead to the loss of subtle subject features, while on the other hand, small cross-attention maps can result in inaccurate optimization between the maps and the downsampled ground truth masks. Regarding the cross-attention maps with middle size from $\mathcal{\ell}_{2}$ and $\mathcal{\ell}_{3}$, $\mathcal{\ell}_{3}$ works better. Thus, in our experiments, we extract cross-attention maps from $\mathcal{\ell}_{3}$ for attention control.
@@ Appendix B Additional Experimental Results
0.6111 0.6335 0.3964 0.7315
0.6080 0.6641 0.3055 0.7242
0.7075 0.6863 0.3983 0.7960
0.6134 0.6420 0.3215 0.7793
0.6234 0.6375 0.3497 0.7443
0.7047 0.6007 0.3642 0.7380
0.6730 0.6364 0.3977 0.7433
0.6508 0.6360 0.3055 0.7469
0.6432 0.6738 0.3594 0.6975
Table 6: Quantitative results of cross-attention levels. We find that only using cross-attention map from $\ell_{3}$ works best and adopt this setting by default.
A fluffy white Persian cat with blue eyes, sitting on a plush red cushion
A playful orange tabby kitten chasing a ball of yarn across a hardwood floor
A sleek black cat with green eyes, perched on a windowsill, looking out at a rainy day
A cute Siamese cat with a pink collar, napping in a wicker basket
A majestic Maine Coon cat with a bushy tail, walking through a field of wildflowers
A curious Sphynx cat with wrinkled skin, exploring a cardboard box
A lazy gray British Shorthair cat, curled up on a soft blanket, basking in the sunlight
A mischievous Calico cat, peeking out from behind a potted plant
A regal Russian Blue cat with silver fur, sitting on a velvet armchair
A friendly Ragdoll cat with bright blue eyes, being petted by a gentle hand
A playful Bengal cat with spots, pouncing on a feather wand toy
A curious Abyssinian cat with large ears, investigating a paper bag
A fluffy Himalayan cat with a flat face, lounging on a fuzzy rug
A sleepy Birman cat with white paws, dozing off on a cozy bed
A graceful Siberian cat with long fur, walking along a wooden fence
A cute Munchkin cat with short legs, playing with a crinkly ball
A curious Norwegian Forest cat with a thick coat, climbing a cat tree
A playful Scottish Fold cat with folded ears, batting at a dangling string
A majestic Savannah cat with a tall, slender build, surveying its surroundings from atop a bookshelf
A friendly Tonkinese cat with a unique coat pattern, rubbing against its owner’s leg
Table 7: Class prompts for generating class images used in class-specific prior preservation. In our experiments, we generate one class image with one independent text prompt. Here we show 20 examples for 20 class images.
Additional Qualitative Results.
We show additional qualitative results here, including comparison of our CustomVideo with previous SOTA methods in Figure 9 and generated video of CustomVideo in Figure 10. It is clear that our method can generate videos with higher fidelity and better preserve the identities of the reference subjects.
Figure 9: Additional comparison results of our CustomVideo with previous SOTA methods.
Figure 10: Additional qualitative results of generated videos from our CustomVideo.
Appendix C Additional Experimental Details
All of the training process is under fp16 mixed precision with accelerate package [62]. We use data augmentations during training, including randomly horizontal flip, randomly crop and resize, together with corresponding prompt change (‘very small’ or ‘close up’ appends before the prompt).
Class Prompts.
For class-specific prior preservation, we generate 200 class images with Stable Diffusion v2.1 [13] for each subject, in which the generating prompts are obtained from Claude-3-Opus [54]. We use Stable Diffusion v2.1 [13] to generate 200 class images for every subject in class-specific prior preservation. The text prompts for generating are obtained from Claude-3-Opus [54]. We show 20 examples in Table 7.
Implementation Details.
We adopt the follow setting for in our experiments, including the previous SOTA methods for comparison and our CustomVideo:
  • DreamBooth [45]: For efficient fine-tuning, we utilize LoRA [50] to adapt the U-Net under DreamBooth [45]. The rank of LoRA module is set as 4. We use AdamW [53] optimizer, with learning rate 4e-4, weight decay 1e-2, batch size 2, and training steps 1000.
  • CustomDiffusion [40]: The weights of query and value in all of the cross-attention layers of U-Net are fine-tuned in CustomDiffusion [40]. We use AdamW [53] optimizer, with learning rate 4e-5, weight decay 1e-2, batch size 2, and training steps 500.
  • VideoDreamer [12]: LoRA [50] is utilized for fine-tuning the U-Net and text encoder in VideoDreamer [12]. The AdamW [53] optimizer is used with learning rate 5e-5, weight decay 1e-2, batch size 2, and training steps 500.
  • DisenDiff + SVD-XT [14, 15]: DisenDiff is trained for 250 steps with AdamW [53] optimizer (learning rate 4e-5, weight decay 1e-2, batch size 2). The sampling step for SVD-XT is 30.
  • DreamVideo [11]: We train the learnable textual token and identity adapter for 3000 and 800 steps, respectively, with learning rate 1e-4 and 1e-5. We use AdamW [53] optimizer with a batch size of 2.
  • CustomVideo: The weights of query and value in all of the cross-attention layers (including the temporal and spatial layers) of U-Net are fine-tuned. The sampling step for high resolution video is 30.
Human Interface for User Study.
The human annotators are all researchers with experience in artificial intelligence. The interface for human evaluation is shown in Figure 11.
Appendix D Broader Impact
Generating realistic videos with human imagination and personal belongings is crucial in practice. Our work makes this poorly explored problem come true by multi-subject driven T2V generation. We believe our work will benefit a wide range of AIGC applications, such as film production and virtual reality. Although our model significantly outperforms other competitive methods, it still struggle to generate promising results for some cases, such as small faces and lots of subjects shown in Figure 7. Please be careful to use our model in the situations where failures will cause severe consequences.
Figure 11: Human interface used in the user study. The generated videos of 5 comparison approaches and our CustomVideo are randomly ordered and anonymous for fair comparison.